OpenStudy (osanseviero):

Tell why this is not a linear tranformation T(a1,a2) = (a1,a1)

OpenStudy (osanseviero):

I checked both properties and both seemed ok.

OpenStudy (osanseviero):

Let's say x=(x1,x2) and y=(y1,y2). I want to prove that T(ax + y) = aT(x) + T(y). T(ax+y) = T(ax1 + y1, ax2 + y2) = (ax1 + y1, ax1 + y1) And aT(x) + T(y) = a(x1,x1)+(y1,y1) = (ax1 + y1, ax1 + y1) They should be different :/

OpenStudy (kainui):

What is \((a_1,a_2)\)? Is it a vector or a matrix with two vectors in it?

OpenStudy (kainui):

I can write this out, \[\begin{pmatrix} 1 & 0 \\ 1 & 0 \end{pmatrix} \begin{pmatrix} a_1 \\ a_2 \end{pmatrix} = \begin{pmatrix} a_1 \\ a_1 \end{pmatrix}\] Looks like it's a linear transformation to me

OpenStudy (osanseviero):

It's a vector with two elements

OpenStudy (osanseviero):

R2->R2

OpenStudy (kainui):

Ok, what's the full question, why do you "know" this isn't a linear transformation? Perhaps they're asking if this is an invertible transformation?

OpenStudy (osanseviero):

The translation is something like this: For the next T: R2 -> R2, explain why it is not linear: a) T(a1,a2)=T(a1,a1)

OpenStudy (kainui):

Ahhh ok this is different

OpenStudy (osanseviero):

Hehe sorry for not being clear.

OpenStudy (kainui):

Unfortunately this is still a linear map whether you write it either way: T(a1,a2)=(a1,a1) or T(a1,a2)=T(a1,a1) :O

OpenStudy (osanseviero):

Weird :S

OpenStudy (osanseviero):

Can you say that T(a1,a2) = T(a1,a1) ??

OpenStudy (kainui):

You said it, not me